# Empowering Time Series Forecasting with LLM-Agents

### v1

@@ 1. Introduction
Time series data pervades diverse real-world applications, including traffic monitoring, financial transactions, and ride-share demand forecasting (Liu et al., 2023). Accurate time series forecasting, in particular, is a critical research area with extensive applications (Godahewa et al., 2021; Zeng et al., 2023; Liu et al., 2023). Concurrently, Large Language Models (LLMs) have demonstrated considerable success in powering Automatic Machine Learning (AutoML) systems for various general machine learning tasks (Huang et al., 2023; Wang et al., 2024a; Schmidt et al., 2024; Chan et al., 2024). This confluence motivates our exploration into developing an LLM-powered AutoML system specifically designed for the unique challenges of time series forecasting.
However, directly applying general AutoML principles to time series forecasting can overlook domain-specific optimization opportunities. Recent findings indicate that lightweight time series models can achieve state-of-the-art performance (Zeng et al., 2023; Chen et al., 2023; Lin et al., 2024; Yeh et al., 2025), shifting focus towards data quality. This aligns with the principles of data-centric AI (Wang et al., 2024b; Zha et al., 2025), which prioritizes data refinement over complex model architectures. We identify a significant opportunity in leveraging LLMs to automate this data-centric approach for time series forecasting, an area that remains underexplored. Therefore, this paper introduces DCATS (Data-Centric Agent for Time Series), an LLM-powered agent that focuses on intelligently refining training data rather than solely optimizing model architectures.
@@
To validate our approach, we conducted a preliminary study employing a large-scale traffic volume forecasting dataset (Liu et al., 2023), notable for its extensive metadata. Our findings demonstrate that DCATS achieves a 6% reduction in forecasting error compared to a baseline model trained using all available time series. This result underscores the efficacy of leveraging LLM-agents for data-centric optimization in time series forecasting and signals a promising avenue for future research across various domains.
@@ 2. Related Work
Traditional AutoML systems primarily focus on model selection, hyperparameter tuning, and pipeline configuration (Thornton et al., 2013; Olson and Moore, 2016; Feurer et al., 2022), employing various optimization and meta-learning techniques to achieve automation. A parallel development has been neural architecture search, which specifically targets the automated design of neural networks using reinforcement learning or evolutionary algorithms (Zoph and Le, 2016; Liu et al., 2018; Pham et al., 2018; Real et al., 2019; Elsken et al., 2019). However, these approaches predominantly emphasize model optimization rather than data quality enhancement. Recent advancements in LLMs have enabled more sophisticated AutoML approaches, with systems like AIDE (Schmidt et al., 2024) employing LLM-agents for solving general machine learning problems. While AIDE demonstrates the potential of LLM-powered AutoML, it does not fully exploit the unique characteristics of time series data. The emerging paradigm of data-centric AI suggests that improving data quality can yield greater performance gains than model refinement alone (Wang et al., 2024b; Zha et al., 2025), particularly relevant for time series forecasting where recent research shows lightweight models with high-quality data can match or exceed complex architectures (Zeng et al., 2023; Chen et al., 2023; Lin et al., 2024; Yeh et al., 2025). Despite these findings, few AutoML systems have been designed with a data-centric focus for time series forecasting, which presents unique challenges including temporal dependencies and domain-specific metadata (Godahewa et al., 2021). To the best of our knowledge, DCATS represents the first LLM-powered data-centric AutoML framework designed specifically for time series forecasting problems, leveraging rich metadata to intelligently select and augment training data.
@@ 3.1. Initial Proposal
Additional guidelines are provided to the agent, as shown in Fig. 5, detailing what is known about each time series and how sub-datasets should be created. Sub-datasets are created by selecting time series from the “neighbors” of the target time series, defined by similarity in road network distance, local correlation, or geodetic distance. These “neighbors” serve a similar function to retrieved documents in a RAG system (Gao et al., 2023; Chang et al., 2024).
@@ 3.2. Forecasting Module
  • Linear (Zeng et al., 2023): A simple yet effective forecasting model.
  • MLP (Chen et al., 2023; Yeh et al., 2024b): A natural extension of the linear model with increased learning capability due to additional parameters.
  • SpraseTSF (Lin et al., 2024): A compact and effective forecasting model.
  • UltraSTF (Yeh et al., 2025): Another compact model with improved cost/learning capability trade-off.
To enhance model convergence speed, we first train a foundation model (Yeh et al., 2023; Miller et al., 2024) using all available time series data before user queries. We then use the sub-dataset proposed by the LLM-agent to fine-tune the foundation model when testing different proposals. To further refine the data before fine-tuning, we remove the 10% most anomalous data from the sub-dataset using discord-based anomaly detection methods (Yeh et al., 2016, 2024a).
@@ 4.1. Dataset and Experiment Setup
We utilized the LargeST dataset (Liu et al., 2023), comprising traffic time series from 8,600 sensors across California. The dataset includes metadata such as location coordinates, county, freeway name, and number of lanes. We augmented this metadata with city names, populations, and historical transaction volumes. Sensor readings were aggregated into 15-minute intervals, resulting in 96 intervals per day over 35,040 time steps. We split each sub-dataset into training, validation, and test sets using a 6:2:2 ratio. Our framework was applied to the training and validation sets, with performance reported on the test set. We tested 60 user queries with randomly selected locations, forecasting the next 12 intervals for each sensor at each timestamp. Performance was evaluated using Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Mean Absolute Percentage Error (MAPE). GPT-4 Turbo was employed as our LLM agent, generating five proposals per round.

### v2

@@ 1. Introduction
Time series data pervades diverse real-world applications, including traffic monitoring, financial transactions, and ride-share demand forecasting (liu2023largest). Accurate time series forecasting, in particular, is a critical research area with extensive applications (godahewa2021monash; zeng2023transformers; liu2023largest). Concurrently, Large Language Models (LLMs) have demonstrated considerable success in powering Automatic Machine Learning (AutoML) systems for various general machine learning tasks (huang2023mlagentbench; wang2024openhands; schmidt2024aide; chan2024mle). This confluence motivates our exploration into developing an LLM-powered AutoML system specifically designed for the unique challenges of time series forecasting.
However, directly applying general AutoML principles to time series forecasting can overlook domain-specific optimization opportunities. Recent findings indicate that lightweight time series models can achieve state-of-the-art performance (zeng2023transformers; chen2023tsmixer; lin2024sparsetsf; yeh2025compact), shifting focus towards data quality. This aligns with the principles of data-centric AI (wang2024visual; zha2025data), which prioritizes data refinement over complex model architectures. We identify a significant opportunity in leveraging LLMs to automate this data-centric approach for time series forecasting, an area that remains underexplored. Therefore, this paper introduces DCATS (Data-Centric Agent for Time Series), an LLM-powered agent that focuses on intelligently refining training data rather than solely optimizing model architectures.
@@
To validate our approach, we conducted a preliminary study employing a large-scale traffic volume forecasting dataset (liu2023largest), notable for its extensive metadata. Our findings demonstrate that DCATS achieves a 6% reduction in forecasting error compared to a baseline model trained using all available time series. This result underscores the efficacy of leveraging LLM-agents for data-centric optimization in time series forecasting and signals a promising avenue for future research across various domains.
@@ 2. Related Work
Traditional AutoML systems primarily focus on model selection, hyperparameter tuning, and pipeline configuration (thornton2013auto; olson2016tpot; feurer2022auto), employing various optimization and meta-learning techniques to achieve automation. A parallel development has been neural architecture search, which specifically targets the automated design of neural networks using reinforcement learning or evolutionary algorithms (zoph2016neural; liu2018darts; pham2018efficient; real2019regularized; elsken2019neural). However, these approaches predominantly emphasize model optimization rather than data quality enhancement. Recent advancements in LLMs have enabled more sophisticated AutoML approaches, with systems like AIDE (schmidt2024aide) employing LLM-agents for solving general machine learning problems. While AIDE demonstrates the potential of LLM-powered AutoML, it does not fully exploit the unique characteristics of time series data. The emerging paradigm of data-centric AI suggests that improving data quality can yield greater performance gains than model refinement alone (wang2024visual; zha2025data), particularly relevant for time series forecasting where recent research shows lightweight models with high-quality data can match or exceed complex architectures (zeng2023transformers; chen2023tsmixer; lin2024sparsetsf; yeh2025compact). Despite these findings, few AutoML systems have been designed with a data-centric focus for time series forecasting, which presents unique challenges including temporal dependencies and domain-specific metadata (godahewa2021monash). To the best of our knowledge, DCATS represents the first LLM-powered data-centric AutoML framework designed specifically for time series forecasting problems, leveraging rich metadata to intelligently select and augment training data.
@@ 3.1. Initial Proposal
Additional guidelines are provided to the agent, as shown in Fig. 5, detailing what is known about each time series and how sub-datasets should be created. Sub-datasets are created by selecting time series from the “neighbors” of the target time series, defined by similarity in road network distance, local correlation, or geodetic distance. These “neighbors” serve a similar function to retrieved documents in a RAG system (gao2023retrieval; chang2024main).
@@ 3.2. Forecasting Module
  • Linear (zeng2023transformers): A simple yet effective forecasting model.
  • MLP (chen2023tsmixer; yeh2024rpmixer): A natural extension of the linear model with increased learning capability due to additional parameters.
  • SpraseTSF (lin2024sparsetsf): A compact and effective forecasting model.
  • UltraSTF (yeh2025compact): Another compact model with improved cost/learning capability trade-off.
To enhance model convergence speed, we first train a foundation model (yeh2023toward; yeh2025tict; yeh2025treasure; miller2024survey) using all available time series data before user queries. We then use the sub-dataset proposed by the LLM-agent to fine-tune the foundation model when testing different proposals. To further refine the data before fine-tuning, we remove the 10% most anomalous data from the sub-dataset using discord-based anomaly detection methods (yeh2016matrix; yeh2024matrix).
@@ 4.1. Dataset and Experiment Setup
We utilized the LargeST dataset (liu2023largest), comprising traffic time series from 8,600 sensors across California. The dataset includes metadata such as location coordinates, county, freeway name, and number of lanes. We augmented this metadata with city names, populations, and historical transaction volumes. Sensor readings were aggregated into 15-minute intervals, resulting in 96 intervals per day over 35,040 time steps. We split each sub-dataset into training, validation, and test sets using a 6:2:2 ratio. Our framework was applied to the training and validation sets, with performance reported on the test set. We tested 60 user queries with randomly selected locations, forecasting the next 12 intervals for each sensor at each timestamp. Performance was evaluated using Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and Mean Absolute Percentage Error (MAPE). GPT-4 Turbo was employed as our LLM agent, generating five proposals per round.